# A Comprehensive Analysis of Social Tie Strength: Definitions, Prediction Methods, and Future Directions

### v2

@@ 5.1 Random Weakly Supervised Setting
  • Nerual network models perform better: As shown in Table 4, the neural network methods outperform heuristic methods, with MLP w/ $\mathbf{E}$ achieving the best performance in the Bitcoin_alpha dataset, and GCN performing the best in CollegeMSG and Twitter datasets. In contrast, heuristic methods depict the worst overall performance in Bitcoin_alpha and CollegeMSG, and the second worst in Twitter. This highlights the greater overall effectiveness of neural network methods in leveraging structural and semantic features to predict tie strength, and heuristic methods that are built on pre-established rules may not be easily adaptable to different real-world scenarios. The fact that MLP w/ $\mathbf{E}$ depicts better than MLP w/o $\mathbf{E}$, but GCN performs better than GTN shows that leveraging semantic features has the potential to improve the performance, but GNNs need better mechanisms to integrate edge features.
@@
  • Models struggle in the most distinct definitions: When analyzing the results across different definitions, we observe that the existing methods vary significantly in their performance. Specifically, Definitions I, V, and VI tend to have lower prediction accuracy. To analyze the reasons for these performance drops, we review the properties related to these definitions and find that these definitions are the most distinct cases according to Figure 2. More specifically, when looking into Definition I, we can find that performance is poor in Bitcoin_alpha and CollegeMSG but better in Twitter. This is likely because in Bitcoin_alpha and CollegeMSG, Definition I is less correlated with other definitions compared to Twitter. A similar pattern is seen in Definitions V and VI. This suggests that 1) there may be limitations in these definitions that cause them to assign pseudo ground-truth labels in a way that differs significantly from other definitions, and 2) model performance is highly dependent on the definition used, and methods struggle when the pseudo ground-truth labels are derived from a perspective different from the common understanding. When definitions reflect the most common understanding of what should be a strong/weak tie, models should be able to perform better.
@@
  • Insiginificant average tie weight differences: The weight of a tie is the sum of all edge weights between its end nodes. Average weight differences between strong/weak ties across different datasets are shown in Table 6. As edge weight reflects the intensity of interactions or the level of trust, we expect the models can successfully reflect this by assigning strong labels to ties with relatively high edge weight. However, compared with the pseudo ground-truth, the existing methods fail to depict significant average weight differences across all definitions. Note that this implies that the existing methods fail to capture the assumed differences in strong versus weak ties having higher and lower edge weight, respectively. This could be because the models suffer from the imbalance issues, and/or are affected by noises in features and local structure information. More discussions are presented in Section 5.4.
5.2 Node-centric Weakly Supervised Setting
@@
The prediction accuracy for node-centric random weakly supervised settings is presented in Table 5. Similar to the results in Section 5.1, neural network methods outperform heuristic methods in Bitcoin_alpha and CollegeMSG, with MLP w/ $\mathbf{E}$ achieving the best performance, while heuristic methods perform poorly in these datasets. However, heuristic methods show the best overall performance on Twitter, mainly due to their strong performance in Definition VI, where the predictors employed by heuristic methods successfully distinguish strong and weak ties. The poorer performance of neural network models here compared to Table 4 is likely because these models rely on hub nodes’ features and local class distribution information, which can introduce bias when this local information deviates significantly from global contexts.
Table 6: Random weakly supervised tie weight difference.
@@
In terms of performance across different definitions, we again observe lower accuracy in Definitions I, V, and VI, consistent with the findings in Section 5.1. This reinforces the observation that model performance depends on the definition used, and models struggle when the labels are derived from a perspective different from the others.
We also notice that models may suffer from imbalance issues, showing biased performance in cases of extreme imbalance, such as Twitter in Definition III and CollegeMSG in Definition IV. Moreover, the average tie weight differences across different datasets (as shown in Table 7) remain less pronounced compared to the pseudo ground-truth. We further discuss these in Section 5.4.
Table 7: Tie weight differences in the Node-centric setting.
@@
(a) Datasets
(b) Definitions
Figure 4: Average weight differences of applying STC-Greedy compared with ground truth across different datasets and definitions. Overall, STC-Greedy depicts low performance and is not applicable for complex situations
5.3 Unsupervised Setting
The inference accuracy of using STC-Greedy is presented in Table 8. Since no label information is provided for training in this setting, the accuracy of STC-Greedy is understandably lower compared to the prediction accuracy discussed in the previous two sections. However, we still observe a notable performance drop in Definitions I, V, and VI, particularly in Bitcoin_alpha and CollegeMSG for Definitions I and V, and in Twitter for Definition VI. This aligns with our previous observations and further demonstrates that model performance is highly dependent on the definition used, with models struggling when pseudo ground-truth labels are derived from different perspectives.
Table 8: Accuracy of STC-Greedy in unsupervised setting.
@@ 5.3 Unsupervised Setting
Table 9: Apply quantity imbalance techniques in the random and node-centric weakly supervised setting.
CollegeMSG IV Twitter III
Methods Acc Macro-F1 Acc Macro-F1
Quantity reweight I
MLP w/o $\mathbf{E}$ 0.853 0.482 0.959 0.492
MLP w/ $\mathbf{E}$ 0.898 0.788 0.971 0.718
GCN 0.908 0.761 0.954 0.504
GTN 0.838 0.502 0.899 0.503
Quantity reweight II
MLP w/o $\mathbf{E}$ 0.878 0.479 0.935 0.492
MLP w/ $\mathbf{E}$ 0.820 0.709 0.967 0.640
GCN 0.859 0.749 0.960 0.491
GTN 0.893 0.473 0.915 0.493
Balanced hub nodes
MLP w/o $\mathbf{E}$ 0.870 0.490 0.953 0.493
MLP w/ $\mathbf{E}$ 0.908 0.798 0.964 0.581
GCN 0.887 0.777 0.959 0.492
GTN 0.825 0.503 0.949 0.496
5.4 Further Investigation
Table 10: Average weight differences between strong and weak ties of MLP w/ $\mathbf{E}$ after applying quantity imbalance techniques is significantly improved.
CollegeMSG IV Twitter III
Ground Truth 14.048 3.143
Quantity reweight I 12.824 2.622
Quantity reweight II 9.587 4.142
Balanced Hub node 14.106 3.169
  • Imbalance Issue: As illustrated in Section 5.1 and  5.2, extreme imbalance cases exist such as CollegeMSG in Definition IV and Twitter in Definition III. To mitigate the imbalance issue, we adopt two quantity imbalance techniques to see if the performance can be improved.
    Quantity reweight (Huang et al. 2016) is one of the well-established quantity imbalance techniques, which assigns larger weight to minority classes during training. In this paper, for edge $e_{ij}\in\mathcal{E}^{Tr}$ that belongs to class $k$ (i.e., $\mathbf{Y}_{ijk}=1,k\in(S,W)$), the quantity reweight $w_{ij}^{q}$ that is applied to it during the training phase is:
    $w_{ij}^{q}=\frac{|\mathcal{E}^{\text{Tr}}|}{|\{\mathcal{E}_{k}^{\text{Tr}}\}_{k=1}^{C}|}$ (1)
    Besides quantity reweight, we also try to mitigate the issues by selecting hub nodes with balanced local class distribution in the node-centric weakly supervised setting.
    The results after applying the quantity imbalance techniques are presented in Table 9, where Quantity reweight I and II denote applying quantity reweight on weakly supervised settings I and II, respectively, and balanced hub nodes represent selecting hub nodes with balanced class distribution in the local neighborhood. We can find even though these quantity imbalance techniques can slightly improve the performance in some cases, i,e., MLP w/ $\mathbf{E}$ for CollegeMSG IV and Twitter III, they all fail to improve the performance accuracy in most other cases significantly. Besides, the Macro-F1 scores remain relatively low, indicating that the models still suffer from imbalance issues. These findings show that the current quantity imbalance techniques may not be sufficient for mitigating the imbalance issues in tie strength prediction tasks. As recent research in imbalanced node/graph classification (Chen et al. 2021; Song, Park, and Yang 2022; Wang et al. 2022) reveals a novel imbalance at the topology level and found its close correlation to the node/graph classification performance, we argue that there could also be a “topological issue” in tie strength prediction that stems from the uneven distributions of local class distributions around edges.
  • Insignificant Average Tie Weight Differences: The imbalance issue highlights that models struggle to differentiate between strong and weak ties. Strong ties typically have larger weights, and this imbalance can contribute to insignificant average weight differences. As demonstrated in Table 10, when applying the two imbalance techniques to MLP w/ $\mathbf{E}$, the average weight differences improve significantly. This suggests that using appropriate imbalance techniques can help to mitigate issues.
    However, these imbalance techniques rely on label information during training, making them unsuitable for unsupervised settings. An alternative improvement strategy is to develop models, such as graph contrastive learning, that can directly separate ties with smaller weights from those with larger weights without requiring label information.
    (a) STC-Greedy
    (b) MLP w/ $\mathbf{E}$
    (c) GCN
    Figure 5: Prediction accuracy of predicting tie strength in the Twitter dataset after $\Delta T$ weeks, based on the initial network.
    (a) Bitcoin_alpha
    (b) CollegeMSG
    (c) Twitter
    Figure 6: Average prediction accuracy of MLP w/ $\mathbf{E}$ and GCN in the random and node-centric weakly supervised setting across different training sample sizes and datasets across 10 runs.
  • Predicting Tie Strength in Dynamic Setting: Building on the work of  (Huang et al. 2018), which shows that tie strength can change over time, we propose that the observed tie strength today may differ from its strength in the future. This suggests that when we predict tie strength based on current data, we might actually be forecasting its future strength rather than its present state. To test this, we ran experiments using three methods (STC-Greedy, MLP w/ $\mathbf{E}$, and GCN) to predict tie strength in the initial Twitter network, comparing the predicted values with the pseudo ground-truth after $\Delta T$ weeks. As shown in Figure 5, the prediction accuracy changes over time for different prediction accuracy. We can observe that Definition VII increases as $\Delta T$ grows, suggesting that in this case, our predictions are more likely forecasting the future strength of a tie. This also implies that in this definition of friend and acquaintance, forming strong friendships takes time, even if conditions for a strong tie are present early on. In contrast, the prediction accuracy for Definition II and III decreases, suggesting that in those cases, predictions are more reflective of the current tie strength rather than its future development, and friendship in these definitions is more likely to be established very soon.
  • Influence of Experiment Setting: To get a deeper understanding of random and node-centric weakly supervised settings and how different models perform under these settings, we vary the size of training samples. The average accuracy of MLP w/ E and GCN on various datasets over 10 runs is shown in Figure 6. Overall, compared to their performance in node-centric supervised settings, models show improved performance in random weakly supervised settings. This might be because the node-centric setting may initially introduce biases based on selected hub nodes’ properties and connectivity patterns, potentially skewing the model’s understanding of the overall network structure. As training sizes increase, however, the gap in performance between the two settings decreases. This observation could indicate that the node-centric setting begins to include a broader variety of edge features and connection patterns. This reduces the initial bias, making the training set more representative of the network as a whole and aligning the performance of the node-centric approach with that of a random edge selection approach.
    Shifting the focus to model comparisons, performance differences between MLP w/ $\mathbf{E}$ and GCN vary across different datasets and training sizes. While GCN tends to outperform MLP w/ $\mathbf{E}$ in larger training scenarios, highlighting its ability to utilize local neighborhood patterns, this advantage is notably significant only in the CollegeMSG dataset. Therefore, further research is needed to understand the conditions under which MLP matches or exceeds GCN performance.
6 Discussion
Based on the previous sections, we conclude the performance of existing models and draw key insights from different perspectives in this section. The limitations and possible future directions are also proposed below:
  • Prediction Model Architectures: Overall, neural network models outperform heuristic methods. This is likely because neural network models are parametrized and can learn the complicated patterns in the data that are slipping from heuristic methods. When training sample size increases, neural network models that can effectively leverage the information in local neighborhoods are more likely to perform better in different scenarios. Besides, semantic features depict the potential to improve performance, but when incorporating them in GNNs, the aggregating and updating mechanism needs careful design. However, despite relatively good performance in weakly supervised settings, neural networks can not be directly implemented for unsupervised tasks as no training labels are provided.
    Also, we can conclude that even though heuristic models with predefined rules might outperform other baselines in biased training scenarios, they generally do not perform well in most other situations. This inconsistency demonstrates that the effectiveness of predefined rules is not guaranteed. For STC-Greedy, as it is not a data-driven learning method, it is hard to perform well in different complex scenarios. Moreover, none of the existing models can distinguish strong and weak ties to reflect the weight differences effectively.
    In this work, we seek to draw valuable insights based on the above observations that help guide the future directions in developing advanced tie strength prediction models. In the future, it would be worthwhile to focus on (1) developing better mechanisms to integrate semantic features into GNN models, (2) extending methods to enable neural networks to perform well even with biased training data by incorporating heuristic rules, and (3) designing graph contrastive learning models (You et al. 2020) that can help distinguish between large and small weight edges for weakly/unsupervised settings.
  • Tie Strength Definitions: Existing models are harder to perform well when pseudo ground-truth labels are derived from a perspective different from the common understanding of strong/weak ties, such as Definition I, V, and VI when the training sample size is relatively small. This also implies that there may be limitations in these definitions. Still, existing models are only good at learning to extract all-rounded information that reflects more of the definitions that align with our common sense. For future research, we can leverage this insight to select the definitions that reflect common understanding and should correspond to the most reasonable definition. Also, we can work on developing models that can perform better even for these distinct definitions, which will increase the models’ performance in different complex scenarios.
  • Weakly Supervised Experiment Setting: Overall, models tend to perform better in the random weakly supervised setting compared to node-centric (as expected), especially with less training data. Additionally, GNN models that effectively use local neighborhood information are more likely to perform well across different settings. One limitation of our study is that we did not thoroughly examine the reasons for performance differences across datasets, e.g., why MLP w/ $\mathbf{E}$ outperforms GCN in Bitcoin_alpha and Twitter but not in CollegeMSG with limited training samples. We believe these differences may stem from the distinct local neighborhood patterns in each dataset and further investigation could lead to models that perform better across a wider range of networks.
  • Tie Strength Imbalance Issues: In many cases, existing models suffer from imbalance issues that we observed cannot be consistently mitigated by quantity imbalance techniques. However, inspired by recent advances in imbalanced edge classification (Cheng et al. 2025), we believe that by examining the imbalance problem from local topological perspectives, an effective solution can be found and one future direction.
  • Predicting Tie Strength in Dynamic Networks: As noted in Section 5.4, different tie strength definitions allow us to infer not only the current tie strength but also its future state. Building on this observation and the insights from tie resilience in Section 3, one of the future research directions to strengthen our current finding could be focusing on developing methods to estimate tie resilience based solely on the current network structure.
7 Conclusions
In this study, we address the challenge posed by the absence of ground-truth tie strength labels and the varying interpretations of tie strength among researchers, which complicate the development of effective prediction methods for real-world applications. To tackle this, we conducted a comprehensive evaluation of existing tie strength prediction methods across a range of real-world definitions. We first categorized mainstream understandings of tie strength into seven standardized definitions and validated their relevance by analyzing class distributions and correlations. Through this analysis, we gained important insights into tie resilience from the tie dissolution perspective. We then conduct extensive experiments and analysis of the performance of existing methods and reveal key findings based on the models’ performance from model architecture and definitions perspectives. We also discovered that current methods suffer from imbalance issues that cannot be adequately addressed by existing quantity imbalance techniques, and different definitions of tie strength enable the prediction of tie strength in different stages of dynamic networks. Based on these, we propose strategies to enhance current models and more importantly, offer several promising directions for future research in this field.

### v5

@@ 7.1. Random Weakly Supervised Setting
  • Nerual network models perform better: As shown in Table 7 (in the Appendix), neural network methods outperform heuristic methods in prediction accuracy, with MLP w/ $\mathbf{E}$ achieving the best performance in the Bitcoin_alpha dataset, and GCN performing the best in CollegeMSG and Twitter datasets. In contrast, heuristic methods depict the worst overall performance in Bitcoin_alpha and CollegeMSG, and the second worst in Twitter. This highlights the greater overall effectiveness of neural network methods in leveraging structural and semantic features to predict tie strength, and heuristic methods that are built on pre-established rules may not be easily adaptable to different real-world scenarios. The fact that MLP w/ $\mathbf{E}$ depicts better than MLP w/o $\mathbf{E}$, but GCN performs better than GTN shows that leveraging semantic features has the potential to improve the performance, but GNNs need better mechanisms to integrate edge features.
@@
  • Models struggle under certain pseudo-labels techniques: When analyzing the results from Figure 4 across different definitions, we observe that the existing methods vary significantly in their performance. Specifically, Definitions I, V, and VI tend to have lower prediction accuracy. To analyze the reasons for these performance drops, we review the properties related to these definitions and find that these definitions are the most distinct cases according to Figure 2. More specifically, when looking into Definition I, we can find that performance is poor in Bitcoin_alpha and CollegeMSG but better in Twitter. This is likely because in Bitcoin_alpha and CollegeMSG, Definition I is less correlated with other definitions compared to Twitter. A similar pattern is seen in Definitions V and VI. This suggests that 1) there may be limitations in these techniques that cause them to assign pseudo ground-truth labels in a way that differs significantly from other definitions, and 2) model performance is highly dependent on the pseudo-labels techniques used. This again highlights the importance of ensuring consistency across studies so that more meaningful comparisons can be made.
@@
  • Insignificant average tie weight differences: The weight of a tie is the sum of all edge weights between its end nodes. Average weight differences between strong/weak ties across different datasets are shown in Table 6. As edge weight reflects the intensity of interactions or the level of trust, we expect the models can successfully reflect this by assigning strong labels to ties with relatively high edge weight. However, compared with the pseudo ground-truth, the existing methods fail to depict significant average weight differences across all definitions, which implies that the existing methods fail to capture the assumed differences in strong versus weak ties having higher and lower edge weights, respectively. This could be because the models suffer from the imbalance issues, and/or are affected by noises in features and local structure information.
Figure 4. Average prediction accuracy of two weakly supervised settings across all models and datasets.
7.2. Node-centric Weakly Supervised Setting
@@
The prediction accuracy for node-centric random weakly supervised settings is presented in Table 8 (in the Appendix). Similar to the results in Section 7.1, neural network methods outperform heuristic methods in Bitcoin_alpha and CollegeMSG, with MLP w/ $\mathbf{E}$ achieving the best performance, while heuristic methods perform poorly in these datasets. However, heuristic methods show the best overall performance on Twitter, mainly due to their strong performance in Definition VI, where the predictors employed by heuristic methods successfully distinguish strong and weak ties. The poorer performance of neural network models here compared to Table 7 is likely because these models rely on hub nodes’ features and local class distribution information, which can introduce bias when this local information deviates significantly from global contexts.
In terms of performance across different definitions, we observe lower accuracy in Definitions I, V, and VI, consistent with Section 7.1. This reinforces that model performance depends on the tie strength definition used; thus, multiple definitions should be used to ensure a robust benchmarking. We also observe that models exhibit biased performance under extreme imbalance, such as Twitter (Definition III) and CollegeMSG (Definition IV). Additionally, average tie weight differences across datasets (Table 9) are less pronounced than those in the pseudo ground-truth; further discussed in Section 7.4. Lastly, we emphasize the need for a node-centric weakly supervised setting beyond the traditional random setting when empirically benchmarking tie strength prediction methods, as it aligns with many real-world scenarios (as discussed in Section ).
(a) Datasets
(b) Definitions
Figure 5. Average weight differences of applying STC-Greedy compared with ground truth across different datasets and definitions. Overall, STC-Greedy depicts low performance and is not applicable for complex situations
7.3. Unsupervised Setting
The inference accuracy of using STC-Greedy is presented in Table 10. Since no label information is provided for training in this setting, the accuracy of STC-Greedy is understandably lower compared to the prediction accuracy discussed in the previous two sections. However, we still observe a notable performance drop in Definitions I, V, and VI, particularly in Bitcoin_alpha and CollegeMSG for Definitions I and V, and in Twitter for Definition VI. This aligns with our previous observations and further demonstrates that model performance is highly dependent on the definition used.
(a) STC-Greedy
(b) MLP w/ $\mathbf{E}$
(c) GCN
Figure 6. Prediction accuracy of predicting tie strength in the Twitter dataset after $\Delta T$ weeks, based on the initial network.
(a) Bitcoin_alpha
(b) CollegeMSG
(c) Twitter
Figure 7. Average prediction accuracy of MLP w/ $\mathbf{E}$ and GCN over all the definitions in the random and node-centric weakly supervised setting across different training sample sizes and datasets across 10 runs.
7.4. Additional Investigation
  • Imbalance Issue: As shown in Section 7.1 and  7.2, extreme class imbalances exist, such as CollegeMSG (Definition IV) and Twitter (Definition III). To address this, we apply two techniques: quantity reweighting and balanced hub node selection.
    Quantity reweighting (Huang et al., 2016) is a common technique that assigns higher weights to minority classes during training. Specifically, for an edge $e_{ij}\in\mathcal{E}^{Tr}$ in class $k$, the assigned weight is:
    (2) $\small w_{ij}^{q}=\frac{|\mathcal{E}^{\text{Tr}}|}{|\{\mathcal{E}_{k}^{\text{Tr}}\}_{k=1}^{C}|}$
    Moreover, in the node-centric setting, we select hub nodes with balanced local class distributions to reduce bias.
    The results in Tables 4 and 5 show that while these techniques improve performance in some cases (e.g., MLP w/ $\mathbf{E}$ for CollegeMSG IV and Twitter III in terms of prediction accuracy), they fail to significantly enhance overall accuracy. Macro-F1 scores remain low, indicating that imbalance remains challenging.
    Recent studies on imbalanced node/graph classification (Chen et al., 2021; Song et al., 2022; Wang et al., 2022b) highlight the impact of topological imbalances on performance. We hypothesize that tie strength prediction also suffers from similar challenges. Inspired by recent imbalance edge classification work (Cheng et al., 2025b), we believe addressing imbalance from a local topological perspective could lead to more effective solutions.
    Table 4. Apply quantity imbalance techniques in the random and node-centric weakly supervised setting.
    CollegeMSG IV Twitter III
    Methods Acc Macro-F1 Acc Macro-F1
    Quantity reweight I
    MLP w/o $\mathbf{E}$ 0.853 0.482 0.959 0.492
    MLP w/ $\mathbf{E}$ 0.898 0.788 0.971 0.718
    GCN 0.908 0.761 0.954 0.504
    GTN 0.838 0.502 0.899 0.503
    Quantity reweight II
    MLP w/o $\mathbf{E}$ 0.878 0.479 0.935 0.492
    MLP w/ $\mathbf{E}$ 0.820 0.709 0.967 0.640
    GCN 0.859 0.749 0.960 0.491
    GTN 0.893 0.473 0.915 0.493
    Balanced hub nodes
    MLP w/o $\mathbf{E}$ 0.870 0.490 0.953 0.493
    MLP w/ $\mathbf{E}$ 0.908 0.798 0.964 0.581
    GCN 0.887 0.777 0.959 0.492
    GTN 0.825 0.503 0.949 0.496
    Table 5. Average weight differences between strong and weak ties of MLP w/ $\mathbf{E}$ after applying quantity imbalance techniques is significantly improved.
    CollegeMSG IV Twitter III
    Ground Truth 14.048 3.143
    Quantity reweight I 12.824 2.622
    Quantity reweight II 9.587 4.142
    Balanced Hub node 14.106 3.169
  • Predicting Tie Strength in a Dynamic Setting: Building on  (Huang et al., 2018), which shows that tie strength evolves, we suggest that today’s observed tie strength may not reflect its future state. This implies that predictions based on current data might forecast future tie strength rather than its present value. To test this, we used three methods (STC-Greedy, MLP w/ $\mathbf{E}$, and GCN) to predict tie strength in the initial Twitter network, then compared the results to the pseudo ground-truth after $\Delta T$ weeks. As shown in Figure 6, how accuracy changes over time depends on the definition used. Definition VII improves as $\Delta T$ increases, suggesting strong friendships take time to form. In contrast, Definitions II and III decline, indicating their predictions reflect current tie strength rather than future development.
  • Impact of Experimental Settings: Based on the performance shown in Figure 4 and in Appendix, overall, existing methods perform similarly in both weakly supervised settings, with the performance on the random setting being slightly better. To explore how models perform under different weakly supervised settings, we varied the training sample size. Figure 7 shows the average accuracy of MLP w/ $\mathbf{E}$ and GCN across datasets over 10 runs. Models generally perform better in the random setting than in the node-centric setting when the training size is small, likely because the latter introduces bias based on hub nodes. However, as training size increases, this gap narrows, suggesting that a broader variety of edges reduces initial bias.
8. Limitations and Future Directions
Since this work presents a comprehensive benchmark for tie strength prediction, our efforts were not to develop the most state-of-the-art prediction method. However, we offer key insights into future advances based on our experimental findings. Overall, neural network models outperform heuristic methods. This is likely because neural network models are parametrized and can learn the complicated patterns in the data that are slipping through heuristic methods. When the training sample size increases, neural network models that can effectively leverage the information in local neighborhoods are more likely to perform better in different scenarios. Besides, semantic features depict the potential to improve performance, but when incorporating them in GNNs, the aggregating and updating mechanism needs careful design. However, despite relatively good performance in weakly supervised settings, neural networks can not be directly implemented for unsupervised tie strength prediction tasks, as no training labels are provided.
Also, we can conclude that even though heuristic models with predefined rules might outperform other baselines in biased training scenarios, they generally do not perform well in most other situations. This inconsistency demonstrates that the effectiveness of predefined rules is not guaranteed. For STC-Greedy, as it is not a data-driven learning method, it is hard to perform well in complex scenarios. Moreover, none of the existing models can distinguish strong and weak ties to reflect the weight differences effectively.
Based on the above, it would be worthwhile to focus on (1) developing better mechanisms to integrate semantic features into GNN models, (2) extending methods to enable neural networks to perform well even with imbalanced training data (Cheng et al., 2025a), i.e., either the inherent label quantity imbalance or structural from the node-centric setting, and (3) designing graph contrastive learning models (You et al., 2020; Wang et al., 2022a) to better leverage unlabeled data that might help distinguish between large and small weight edges for weakly/unsupervised settings.
9. Conclusion
In this study, we address the challenge in current tie strength prediction research by introducing BTS, a comprehensive Benchmark for Tie Strength prediction. Specifically, we categorize mainstream understandings of tie strength into seven standardized definitions and validate their effectiveness by analyzing class distributions and correlations across these definitions on three well-curated real-world datasets. Then we propose a standardized framework to evaluate the quality of pseudo-labels from the perspective of tie resilience. To benchmark tie strength prediction, we evaluate the performance of representative models, and explore the effects of different experimental settings, models, and evaluation criteria on prediction performance. Overall, BTS establishes a standardized foundation for evaluating and advancing tie strength prediction methodologies.
10. Acknowledgments
This research is supported by the National Science Foundation under grant number IIS2239881.
Appendix A Appendix
Below, we present all additional experimental results, including prediction accuracy and tie weight differences for both weakly supervised settings, as well as STC-Greedy performance under the unsupervised setting.
Table 6. Tie weight difference (random weakly setting).
@@ Appendix A Appendix
Table 7. Prediction accuracy of existing models on three datasets across seven definitions in the random weakly supervised setting. Overall, models perform worse in Definition I, V, and VII, with MLP w/ $\mathbf{E}$ and GCN performing best.
Definitions Model Bitcoin_alpha CollegeMSG Twitter Definitions Model Bitcoin_alpha CollegeMSG Twitter
MLP w/o $\mathbf{E}$ 0.653 0.532 0.756 MLP w/o $\mathbf{E}$ 0.521 0.610 0.731
MLP w/ $\mathbf{E}$ 0.789 0.845 0.909 MLP w/ $\mathbf{E}$ 0.560 0.845 0.743
GCN 0.719 0.824 0.885 GCN 0.537 0.809 0.713
GTN 0.640 0.552 0.695 GTN 0.507 0.517 0.624
MLP w/o $\mathbf{E}$ 0.794 0.758 0.930 MLP w/o $\mathbf{E}$ 0.741 0.781 0.573
MLP w/ $\mathbf{E}$ 0.783 1.000 0.939 MLP w/ $\mathbf{E}$ 0.726 0.900 0.544
GCN 0.794 0.976 0.931 GCN 0.674 0.863 0.637
GTN 0.789 0.757 0.931 GTN 0.650 0.794 0.558
MLP w/o $\mathbf{E}$ 0.828 0.789 0.865 MLP w/o $\mathbf{E}$ 0.766 0.830 0.865
MLP w/ $\mathbf{E}$ 0.816 0.965 0.867 MLP w/ $\mathbf{E}$ 0.756 0.907 0.867
GCN 0.828 0.939 0.960 GCN 0.731 0.907 0.867
GTN 0.825 0.789 0.960 GTN 0.666 0.830 0.841
MLP w/o $\mathbf{E}$ 0.889 0.882 0.843 MLP w/o $\mathbf{E}$ 0.742 0.740 0.808
MLP w/ $\mathbf{E}$ 0.882 0.890 0.845 MLP w/ $\mathbf{E}$ 0.759 0.907 0.830
GCN 0.889 0.905 0.845 GCN 0.739 0.889 0.834
GTN 0.886 0.894 0.845 GTN 0.709 0.733 0.779
Table 8. Prediction accuracy of existing models on three datasets across seven definitions in the node-centric weakly supervised setting. Similarly, overall models perform worse in Definition I, V, and VII, and MLP w/ $\mathbf{E}$ and GCN perform best.
MLP w/o $\mathbf{E}$ 0.746 0.506 0.663 MLP w/o $\mathbf{E}$ 0.526 0.638 0.735
MLP w/ $\mathbf{E}$ 0.833 0.817 0.914 MLP w/ $\mathbf{E}$ 0.533 0.813 0.736
GCN 0.741 0.817 0.757 GCN 0.516 0.798 0.733
GTN 0.650 0.511 0.743 GTN 0.495 0.588 0.685
MLP w/o $\mathbf{E}$ 0.794 0.757 0.931 MLP w/o $\mathbf{E}$ 0.741 0.794 0.454
MLP w/ $\mathbf{E}$ 0.794 0.918 0.931 MLP w/ $\mathbf{E}$ 0.729 0.877 0.458
GCN 0.794 0.884 0.931 GCN 0.741 0.855 0.614
GTN 0.760 0.758 0.931 GTN 0.712 0.776 0.524
MLP w/o $\mathbf{E}$ 0.828 0.789 0.960 MLP w/o $\mathbf{E}$ 0.766 0.829 0.867
MLP w/ $\mathbf{E}$ 0.828 0.969 0.960 MLP w/ $\mathbf{E}$ 0.766 0.883 0.867
GCN 0.828 0.885 0.960 GCN 0.766 0.832 0.853
GTN 0.805 0.789 0.960 GTN 0.736 0.830 0.853
MLP w/o $\mathbf{E}$ 0.889 0.894 0.845 MLP w/o $\mathbf{E}$ 0.756 0.744 0.779
MLP w/ $\mathbf{E}$ 0.889 0.894 0.845 MLP w/ $\mathbf{E}$ 0.767 0.882 0.816
GCN 0.889 0.894 0.845 GCN 0.754 0.852 0.813
GTN 0.882 0.894 0.803 GTN 0.720 0.735 0.786
Table 9. Tie weight difference (node-centric setting).
@@
Table 10. Accuracy of STC-Greedy in unsupervised setting.